solve the following system of equations using the elimination method:
3x+2y=48
9x-8y=-24

Answer by Maths68(1474)

3x+2y=48........(1)
9x-8y=-24........(2)

Multiply (1) by 3 and subtract it from (2)

9x-8y=-24
-9x-6y=-144
-14y=-168
-14y/-14=-168/-14
y=12
Put the value of y in (1)
3x+2y=48
3x+2(12)=48
3x+24=48
3x=48-24
3x=24
3x/3=24/3
x=8
x,y=8,12
Check
=====
Put the values of x and y in (2)
9x-8y=-24
9(8)-8(12)=-24
72-96=-24
-24=-24